Demand for a country's currency in the foreign exchange market is given by where XR is...

  • Demand for a country's currency in the foreign exchange market is given byLaTeX: XR\:=\:200\:-0.5A where XR is the US dollar price of the currency, and A is the quantity of the currency.
  • Supply for Country A's currency in the foreign exchange market is given by LaTeX: XR=10+0.5A
  • The central bank of the country fixes the exchange rate at 62 USD.

The central bank needs to sell A = ________ in the foreign exchange market to maintain the fixed exchange rate.

[Fill in the blank with ONLY numerical values. For example, if the central bank needs to sell 100 units of the domestic currency, enter 100.]
Answer #1

The Demand of country's currency = XR=200-0.5A

where,A is quantity of currency,XR is US Dollar price.

The central bank wants to the exchange rate at 62 USD

Lets calculate the demand for currency

XR=200-0.5A

62=200-0.5A

0.5A= 200-62

A= 138/0.5=276

Supply of country's currency

XR=10+0.5A

62=10+0.5A

62-10=0.5A

A= 52/0.5=104

Hence Demand is more and supply of the currency is less. So the price of USD is likely to increase

So in order to manitain the exchange rate at 62 USD

The cenral bank needs to sell A=(276-104) 172 in order to manitain the fixed exchange rate.

So that supply icreases and exchange rate reamins constant.